# Anticipatory Music Transformer

### v1

@@ 3.4 Anticipatory Inference
We draw conditional samples from an anticipatory autoregressive model $p$ according to the procedure described in Algorithm 1. If there are controls ($K\neq 0)$ we set $\mathbf{z}=1$ (anticipatory sampling mode). In the outer loop, we sample $\mathbf{a}_{i}\sim p(\cdot|\mathbf{a}_{1:i-1})$. Each time we draw a sample $\mathbf{a}_{i}$ from the model, we note its time $t_{i}=\texttt{time}(\mathbf{a}_{i})$ and check for controls satisfying the condition given by Equation (4) (leveraging the fact that this condition is a stopping time). We anticipate these controls by appending them to the generated output following $\mathbf{a}_{i}$. The sequence $\mathbf{a}_{1:N+K}$ returned by Algorithm 1 contains $N$ events interleaved with $K$ labels. We can postprocess this sequence by stripping out the controls to recover the sequence of generated events $\mathbf{e}_{1:N}$.
@@ 3.5 Anticipatory Infilling Models
We describe the precise encoding of $\mathcal{U}\cup\mathcal{V}$ that we use for music infilling models in Appendix B.
Previously we have assumed that the controls $\mathbf{u}_{1:K}$ are given to us, separate from the the events $\mathbf{e}_{1:N}$. For the infilling task, there is no distinction between events and controls in the training data; a user could specify an arbitrary set of events to condition on at inference time. To train an infilling model, we must impose a prior over the distribution over subsets of control events $\mathbf{u}_{1:K}\subseteq\mathbf{e}_{1:N}$. This prior should simulate common infilling patterns, and generalize to accommodate patterns that we did not foresee during training. To that end, we propose a distribution over control events consisting of a mixture of random spans of time, random subgroups of instrument, and uniformly random events. We describe specific details of this prior distribution in Appendix C.
We caution that the conditionals learned by an anticipatory infilling model are not consistent with a unique joint distribution over sequences $\mathbf{e}_{1:N}$. Given controls $\mathbf{u}_{1:K}$, the map between the sequences $\mathbf{e}_{1:N}$ and $\mathbf{a}_{1:N}$ (Equation (17)) is a bijection. The probability distribution over sequences $\mathbf{a}_{1:N}$ therefore also defines a distribution over sequences $\mathbf{e}_{1:N}$. Every subset $\varphi^{\raisebox{0.60275pt}{$\scriptscriptstyle-1$}}(\mathbf{u}_{1:K})\subseteq\mathbf{e}_{1:N}$ can be used as infilling controls, each resulting in a distinct anticipatory sequence $\mathbf{a}_{1:N}$. But the probability distributions $p(\mathbf{e}_{1:N})$ implied by sequences $\mathbf{a}_{1:N}$ will not be the same. We rely on the model’s ability to well-approximate the underlying distribution over sequences $\mathbf{e}_{1:N}$ to ensure some degree of compatibility between these families of learned distributions.
@@ 4 Anticipatory Infilling Models of Music
We train and release anticipatory infilling models on the Lakh MIDI dataset (Raffel, 2016). See Appendix A for licensing information and Section 7 for a discussion of copyright considerations regarding models trained on Lakh MIDI. The Lakh MIDI dataset consists of $178,561$ MIDI files (event sequences) that we preprocess into $663,555,310$ events ($1,990,665,930$ tokens using arrival-time encoding) encompassing $8943$ hours of music. This dataset is orders of magnitude larger than other common music datasets (Dong et al., 2020), but orders of magnitude smaller than the datasets used to train large models in other domains, e.g. language (Gao et al., 2020) and vision (Schuhmann et al., 2022). For reference, the OpenWebText corpus that reproduces the training set used to train GPT-2 contains approximately $10$ billion tokens (Gokaslan and Cohen, 2019). We describe training splits and additional information about Lakh MIDI and preprocessing in Appendix D.
We parameterize anticipatory infilling models using causal masked transformers (Vaswani et al., 2017) (decoder-only models) with anticipation interval $\delta=5$ seconds. This allows the models to look ahead (i.e. anticipate) controls up to $5$ seconds before their arrival. Due to the high cost of training, we defer systematic ablations of this hyper-parameter to future work. We train models at three scales, following GPT-2 naming conventions (Radford et al., 2019): Small ($128$M parameters), Medium ($360$M parameters), and Large ($780$M parameters). Because anticipatory models are trained like sequence models on the augmented sequences $\mathbf{a}_{1:N+K}$ (see Section 3.3) we are able to use standard libraries for training anticipatory music transformers; in this work, we use the Levanter library for training. [^2] https://github.com/stanford-crfm/levanter For additional details of models and training procedures, see Appendix F.
@@ 4.1 Automatic Metrics
Table 1: Evaluation Loss. All losses are reported on the Lakh MIDI test set sequences $\mathbf{e}_{1:N}$ (without anticipation). For the arrival-time models, we define a per-event loss $L_{\mathbf{e}}$ summed across event triples (Definition 2.2) and report the per-event perplexity $\text{ppl}(\mathbf{e})=\exp\left(L_{\mathbf{e}}\right)$. This loss decomposes into $L_{\mathbf{e}}=L_{\mathbf{t}}+L_{\mathbf{d}}+L_{\mathbf{n}}$ with corresponding perplexities for onsets $\mathbf{t}$, durations $\mathbf{d}$ and notes $\mathbf{n}$. The time-normalized bits per second metric (bps) (Thickstun et al., 2019) is defined in Appendix E. Parameter counts differ between arrival and interarrival models, due to the difference in vocabulary size. Anticipatory models are indicated by the ‘AM’ flag. For anticipatory models, we evaluate with $\textbf{z}=0$ (see Section 3.3).
@@
See Table 1 for a summary of configurations and performance metrics derived from the log-loss for anticipatory infilling models trained on Lakh MIDI. Arrival-time encoding appears more effective for training autoregressive transformers than interarrival-time encoding (rows $1$ and $2$). Anticipatory training incurs a small tax on performance compared to a baseline autoregressive model for short training schedules (compare rows 2 and 3) however this gap disappears with a longer training schedule (compare rows 4 and 5). In this sense, anticipation unlocks infilling control capabilities in autoregressive music models “for free.” Larger models achieve significantly better perplexity on test (compare rows 3, 6, and 9) as do longer training schedules (compare rows 2 and 4; 3 and 5; 6, 7, and 8). Therefore, despite the relatively small size of the Lakh MIDI dataset compared to language modeling datasets, we do not appear to have saturated this dataset or overfit: better results are likely achievable with larger-scale models, trained for longer.
@@ 4.2 Human Evaluation
Workers evaluated $50$ clips, each consisting of a three-bar prompt—randomly selected from the beginning of tracks in the Lakh MIDI test set—followed by a continuation generated by a model or from the original human composition. See Table 2 for human evaluation of music generated from prompts, compared to human compositions; for full pairwise comparison results, see Appendix I. We use this task to compare our model in the standard autoregressive setting to a baseline Music Transformer (Huang et al., 2018) as implemented by von Rütte et al. (2023) and trained on the Lakh MIDI dataset; we refer to this as the FIGARO baseline. While continuing a prompt does not require anticipation, evaluators find that music generated by anticipatory models given a prompt is considerably more musical than music generated by the baseline FIGARO model (this should not be surprising; the FIGARO Music Transformer is a smaller, $30$M parameter model). Human evaluators also prefer the Small arrival-time model (Row 3; Table 1) over the Small interarrival-time model (Row 1; Table 1) which is consistent with the superior log-loss of the arrival-time model.
@@
Workers evaluated $20$ clips, each consisting of a five-second prompt (randomly selected from tracks in the Lakh MIDI test set) followed by a $15$-second accompaniment (i.e. infilling) conditioned on the prompt and the full ($20$ seconds) melodic line. See Table 3 for human evaluation of generated music accompaniments, compared to human compositions; for full pairwise comparison results, see Appendix I. For this task, we crudely define melodic line to be the instrumental part with the highest pitch. This task allows us to probe the infilling capabilities of our anticipatory autoregressive models. We compare anticipation to two baseline accompaniment algorithms: random retrieval and autoregressive accompaniment. Random retrieval is a simple baseline whereby we accompany the melody with a random $15$-second clip retrieved from elsewhere in the track. Autoregressive accompaniment is a best effort algorithmic attempt to use autoregressive generation (without anticipation) to solve the accompaniment task; see Appendix H for details. Both anticipatory and autoregressive accompaniments are generated using a Medium anticipatory model (Row 8; Table 1). Evaluators express a mild preference for anticipatory accompaniments over the human composition skyline. While this result is not statistically significant, it points to the effectiveness of anticipation, especially in the more constrained accompaniment setting.
Related Work on Musical Accompaniment.
In contrast to harmonization tasks, which seek to accompany a melody with simple chords (Simon et al., 2008; Yeh et al., 2021; Chen et al., 2021; Wu et al., 2021; Rhyu et al., 2022), here we seek to generate complete, asynchronous musical accompaniments to a given melody. Dong et al. (2018) previously considered an accompaniment task using piano-roll encodings of the Lakh MIDI dataset: see Section 2 for a discussion of the limitations of piano-roll encodings. Both Zhu et al. (2018) and Ren et al. (2020) propose encoder-decoder architectures for conditional generation of accompaniments given a melody: we discuss obstructions to efficiently training encoder-decoder models of music at scale in Section 1. Shih et al. (2022) consider a more abstract form of conditioning, using a melody as thematic material to generate an arrangement—rather than a literal accompaniment—of the given melody. Our accompaniment task is also loosely analogous to recent work on vocal accompaniment in the audio domain (Donahue et al., 2023).
@@
5 Related Work on Controllable Generative Modeling
Markov-chain Monte Carlo (MCMC) samplers are a natural candidate for conditional sampling tasks, and in particular infilling. Diffusion models have proven highly effective as generative models of images with infilling capabilities (Ho et al., 2020; Lugmayr et al., 2022). Masked language models can also be repurposed as MCMC samplers for infilling (Goyal et al., 2021; Mireshghallah et al., 2022; Wang and Xia, 2021). These models are not easily applied to conditional point process generation, which requires the generation of variable numbers of events within a specified region of space or time. MCMC models of symbolic music compromise by either discretizing time with piano-rolls (Hadjeres et al., 2017; Huang et al., 2017) or modeling coarse, fixed-rate latent variables, without the ability to control granular details of the variable-length sequence (Mittal et al., 2021).
@@ 6 Conclusion
Provided that locality continues to be an important inductive bias for learning, we believe that anticipation will be useful for modeling conditional temporal point processes. Locality is a perennial theme in the machine learning literature. Popular model architectures including ConvNets (Forsyth et al., 1999) and LSTM (Hochreiter and Schmidhuber, 1997) exploit locality as an inductive bias. Many long context Transformers adopt local sparsity as an approximation to dense attention (Child et al., 2019; Beltagy et al., 2020; Zaheer et al., 2020). Locality can be exploited to improve training efficiency via staged training (Press et al., 2021, 2022). The local structure of anticipatory sequences is conducive to the broad class of methods that exploit or depend upon locality. Because anticipation only intervenes to modify the data, it can be seamlessly combined with other modeling innovations, e.g., in the music space: the relative transformer (Huang et al., 2018), the compound word transformer (Hsiao et al., 2021), and RIPO attention (Guo et al., 2022).
@@
Generative music models have not yet reached a broad audience alongside generative models of language (Brown et al., 2020) or images (Ramesh et al., 2022). Slow adoption of music models within creative communities is partly attributable to the difficulty of controlling these models (Briot and Pachet, 2020). Users value agency in human-AI collaborations, preferring to take an active role over more automated solutions (Oh et al., 2018; Roy et al., 2019). We note the proliferation of recent work on music generation (primarily in the audio domain) controlled by text (Forsgren and Martiros, 2022; Kreuk et al., 2022; Agostinelli et al., 2023; Schneider et al., 2023; Huang et al., 2023a, b); we are intrigued by the possibility of applying anticipation to generate symbolic music conditioned on localized text labels (e.g., lyrics). More broadly, we view symbolic music generation and audio music synthesis as complementary, analogous to text generation and speech synthesis. We hope that the Anticipatory Music Transformer may fill a role that is currently underdeveloped in the construction of controllable music generation systems that support the human creative process.
7 Ethical Considerations for Generative Models of Music

### v2

@@ 3.4 Anticipatory Inference
We draw conditional samples from an anticipatory autoregressive model $p$ according to the procedure described in Algorithm 1. If there are controls ($K\neq 0)$ we set $\mathbf{z}=1$ (anticipatory sampling mode). In the outer loop, we sample $\mathbf{a}_{i}\sim p(\cdot|\mathbf{a}_{1:i-1})$. We impose three constraints upon the sampling distribution, all implemented by logit masking: (1) each arrival time must equal or exceed the previous arrival time (monotonicity) (2) a time token must follow a note token, a duration token must follow a time token, and a note token must follow a duration token (proper ordering) and (3) the model must not generate a control token (controls are pre-specified by the user).
Each time we draw a sample $\mathbf{a}_{i}$ from the model, we note its time $t_{i}=\texttt{time}(\mathbf{a}_{i})$ and check for controls satisfying the condition given by Equation (4) (leveraging the fact that this condition is a stopping time). We anticipate these controls by appending them to the generated output following $\mathbf{a}_{i}$. The sequence $\mathbf{a}_{1:N+K}$ returned by Algorithm 1 contains $N$ events interleaved with $K$ labels. We can postprocess this sequence by stripping out the controls to recover the sequence of generated events $\mathbf{e}_{1:N}$.
@@ 3.5 Anticipatory Infilling Models
We describe the precise encoding of $\mathcal{U}\cup\mathcal{V}$ that we use for music infilling models in Appendix C.
Previously we assumed that the controls $\mathbf{u}_{1:K}$ are given to us, separate from the the events $\mathbf{e}_{1:N}$. For the infilling task, there is no distinction between events and controls in the training data; a user could specify an arbitrary set of events to condition on at inference time. To train an infilling model, we must impose a prior on the distribution over subsets of control events $\mathbf{u}_{1:K}\subseteq\mathbf{e}_{1:N}$. This prior should simulate common infilling patterns, and generalize to accommodate patterns we did not foresee during training. To that end, we propose a distribution over control events consisting of a mixture of random spans of time, random subgroups of instrument, and uniformly random events. We describe specifics of this prior in Appendix D.
We caution that the conditionals learned by an anticipatory infilling model are not consistent with a unique joint distribution over sequences $\mathbf{e}_{1:N}$. Given controls $\mathbf{u}_{1:K}$, the map between the sequences $\mathbf{e}_{1:N}$ and $\mathbf{a}_{1:N}$ defined by Equation (17) is a bijection. The probability distribution over sequences $\mathbf{a}_{1:N}$ therefore also defines a distribution over sequences $\mathbf{e}_{1:N}$. Every subset $\varphi^{\raisebox{0.60275pt}{$\scriptscriptstyle-1$}}(\mathbf{u}_{1:K})\subseteq\mathbf{e}_{1:N}$ can be used as infilling controls, each resulting in a distinct anticipatory sequence $\mathbf{a}_{1:N}$. But the probability distributions $p(\mathbf{e}_{1:N})$ implied by sequences $\mathbf{a}_{1:N}$ will not be the same. We rely on the model’s ability to well-approximate the data distribution to ensure approximate consistency between these families of learned distributions.
@@ 4 Anticipatory Infilling Models of Music
We train and release anticipatory infilling models on the Lakh MIDI dataset (Raffel, 2016). See Appendix B for licensing information and Section A for a discussion of copyright considerations regarding models trained on Lakh MIDI. The Lakh MIDI dataset consists of $178,561$ MIDI files (event sequences) that we preprocess into $663,555,310$ events ($1,990,665,930$ tokens using arrival-time encoding) encompassing $8943$ hours of music. This dataset is orders of magnitude larger than other common music datasets (Dong et al., 2020), but orders of magnitude smaller than the datasets used to train large models in other domains, e.g. language (Gao et al., 2020) and vision (Schuhmann et al., 2022). For reference, the OpenWebText corpus that reproduces the training set used to train GPT-2 contains approximately $10$ billion tokens (Gokaslan & Cohen, 2019). We describe training splits and additional information about Lakh MIDI and preprocessing in Appendix E.
All models trained in this paper are parameterized using causal masked transformers (Vaswani et al., 2017) (decoder-only models) with a context length of $M=$ $1024$ tokens (defined in Section 3.3). We train models at three scales, following GPT-2 naming conventions (Radford et al., 2019): Small ($128$M parameters), Medium ($360$M parameters), and Large ($780$M parameters). Because anticipatory models are trained like sequence models on the augmented sequences $\mathbf{a}_{1:N+K}$ (see Section 3.3) we are able to use standard libraries for training anticipatory music transformers; in this work, we use the Levanter library for training. [^2] https://github.com/stanford-crfm/levanter For additional details of models and training procedures, see Appendix G.
Anticipatory models are trained with an anticipation interval of $\delta=$ $5$ seconds; this allows the models to look ahead (i.e. anticipate) controls up to $5$ seconds before their arrival. This interval is chosen to maximize anticipation, while not anticipating controls so far in advance that a control on time $s$ might appear more than $M$ tokens earlier than events near time $s$ (outside of the model context $M$). Music in the Lakh MIDI dataset is represented using an average of $68$ tokens per second, with a standard deviation of $51$ tokens per second (see Appendix E). Therefore, setting $\delta=$ $5$ seconds ensures that controls on time $s$ appear within $M=$ $1024$ tokens of events near time $s$, unless tokens are generated at a sustained rate of more than 2.67 standard deviations above the mean (i.e., more than $204\approx 68+2.67*51$ tokens/second $\approx M/\delta$).
@@ 4.1 Automatic Metrics
Table 1: Evaluation Loss. All losses are reported on the Lakh MIDI test set sequences $\mathbf{e}_{1:N}$ (without anticipation). For the arrival-time models, we define a per-event loss $L_{\mathbf{e}}$ summed across event triples (Definition 2.2) and report the per-event perplexity $\text{ppl}(\mathbf{e})=\exp\left(L_{\mathbf{e}}\right)$. This loss decomposes into $L_{\mathbf{e}}=L_{\mathbf{t}}+L_{\mathbf{d}}+L_{\mathbf{n}}$ with corresponding perplexities for onsets $\mathbf{t}$, durations $\mathbf{d}$ and notes $\mathbf{n}$. The time-normalized bits per second metric (bps) (Thickstun et al., 2019) is defined in Appendix F. Parameter counts differ between arrival and interarrival models, due to the difference in vocabulary size. Anticipatory models are indicated by the ‘AM’ flag. For anticipatory models, we evaluate with $\textbf{z}=0$ (see Section 3.3).
@@
See Table 1 for a summary of configurations and performance metrics derived from the log-loss for anticipatory infilling models trained on Lakh MIDI. Arrival-time encoding appears more effective for training autoregressive transformers than interarrival-time encoding (rows $1$ and $2$). Anticipatory training incurs a small tax on performance compared to a baseline autoregressive model for short training schedules (compare rows 2 and 3) however this gap disappears with a longer training schedule (compare rows 4 and 5). In this sense, anticipation unlocks infilling control capabilities in autoregressive music models “for free.” Larger models achieve significantly better perplexity (compare rows 3, 6, and 9) as do longer training schedules (compare rows 2 and 4; 3 and 5; 6, 7, and 8).
@@ 4.2 Human Evaluation
Workers evaluated $50$ clips, each consisting of a three-bar prompt—randomly selected from the beginning of tracks in the Lakh MIDI test set—followed by a continuation generated by a model or from the original human composition. See Table 2 for human evaluation of music generated from prompts, compared to human compositions; for full pairwise comparison results, see Appendix J. We use this task to compare our model in the standard autoregressive setting to a baseline Music Transformer (Huang et al., 2018) as implemented by von Rütte et al. (2023) and trained on the Lakh MIDI dataset; we refer to this as the FIGARO baseline. While continuing a prompt does not require anticipation, evaluators find that music generated by anticipatory models given a prompt is considerably more musical than music generated by the baseline FIGARO model (this should not be surprising; the FIGARO Music Transformer is a smaller, $30$M parameter model). Human evaluators prefer the Small arrival-time model (Row 3; Table 1) over the Small interarrival-time model (Row 1; Table 1) consistent with the superior log-loss of the arrival-time model.
@@
Workers evaluated $20$ clips, each consisting of a five-second prompt (randomly selected from tracks in the Lakh MIDI test set) followed by a $15$-second accompaniment (i.e. infilling) conditioned on the prompt and the full ($20$ seconds) melodic line. See Table 3 for human evaluation of generated music accompaniments, compared to human compositions; for full pairwise comparison results, see Appendix J. For this task, we crudely define melodic line to be the instrumental part with the highest pitch.
This task allows us to probe the infilling capabilities of our anticipatory autoregressive models. We compare anticipation to two baseline accompaniment algorithms: random retrieval and autoregressive accompaniment. Random retrieval is a simple baseline whereby we accompany the melody with a random $15$-second clip retrieved from elsewhere in the track. Autoregressive accompaniment is an algorithmic attempt to use autoregressive generation (without anticipation) to solve the accompaniment task; see Appendix I for details. Both anticipatory and autoregressive accompaniments are generated using a Medium anticipatory model (Row 8; Table 1). Evaluators express a mild preference for anticipatory accompaniments over the human composition skyline. While this result is not statistically significant, it points to the effectiveness of anticipation, especially in the more constrained accompaniment setting.
@@
5 Related Work
Controllable Generative Modeling.
@@ 5 Related Work
Markov-chain Monte Carlo (MCMC) samplers are a natural candidate for conditional sampling tasks, and in particular infilling. Diffusion models have proven highly effective as generative models of images with infilling capabilities (Ho et al., 2020; Lugmayr et al., 2022). Masked language models can also be repurposed as MCMC samplers for infilling (Goyal et al., 2021; Mireshghallah et al., 2022; Wang & Xia, 2021). These models are not easily applied to conditional point process generation, which requires the generation of variable numbers of events within a specified region of space or time; for work developing MCMC infilling methods in restricted families of point processes, see Shelton et al. (2018). Applications of MCMC methods for symbolic music infilling have compromised by either discretizing time with piano-rolls (Hadjeres et al., 2017; Huang et al., 2017) or modeling coarse, fixed-rate latent variables, without the ability to control granular details of the variable-length sequence (Mittal et al., 2021).
Sequential Monte Carlo (SMC) samplers have been considered for infilling of both continuous temporal point processes (Mei et al., 2019) and sequence models (Lin & Eisner, 2018); the later more directly relates to our setting because discretizing time (Section 2.1) reduces point process modeling to sequence modeling. These SMC methods approximate samples from the conditional distribution over events given all past and future controls, using importance sampling to select among candidate samples offered by a proposal model. In contrast to SMC, anticipatory inference is more efficient, requiring just one model call to generate an event. The limitations of anticipation are also more explicit: conditional independence of controls more than $\delta$ seconds in the future. In contrast, the limitations of SMC approximations are implicitly determined by the number of candidates and the effectiveness of the proposal model.
Musical Accompaniment.
In contrast to harmonization tasks, which seek to accompany a melody with simple chords (Simon et al., 2008; Yeh et al., 2021; Chen et al., 2021; Wu et al., 2024; Rhyu et al., 2022), here we seek to generate complete, asynchronous musical accompaniments to a given melody. Huang et al. (2018) previously considered a solo-piano accompaniment generation task using Seq2Seq models; anticipatory models generalize the Seq2Seq approach, motivated by the long multi-instrument sequences found in the Lakh MIDI dataset. Dong et al. (2018) considered an accompaniment task using piano-roll encodings of the Lakh MIDI dataset; see Section 2 for a discussion of the limitations of piano-roll encodings. Both Zhu et al. (2018) and Ren et al. (2020) propose encoder-decoder architectures for conditional generation of accompaniments given a melody: we discuss obstructions to efficiently training encoder-decoder models of music at scale in Section 1. Shih et al. (2022) consider a more abstract form of conditioning, using a melody as thematic material to generate an arrangement—rather than a literal accompaniment—of the given melody. Our accompaniment task is also loosely analogous to recent work on vocal accompaniment in the audio domain (Donahue et al., 2023).
@@ 6 Conclusion
While the focus of this work has been the music domain, the anticipatory modeling techniques developed in Section 3 are generally applicable to the controllable generation of temporal point processes. Temporal point processes appear in any setting where data is associated with timestamps, including ecommerce activity (Du et al., 2015), social media logs (Farajtabar et al., 2017), healthcare records (Enguehard et al., 2020), and neuroscience data (Williams et al., 2020). Using analogs to the encoding developed for music in Section 2, anticipatory modeling could be used to facilitate the controllable simulation of data in these and other modalities. Furthermore, while the experiments in this paper focus on infilling control–a special case where the controls and events belong to the same modality–anticipatory modeling could be applied more generally to simulate events in one modality, guided asynchronous by controls in a different modality.
@@
Provided that locality continues to be an important inductive bias for learning, we believe that anticipation will be useful for modeling conditional temporal point processes. Locality is a perennial theme in the machine learning literature. Popular model architectures including ConvNets (Forsyth et al., 1999) and LSTM (Hochreiter & Schmidhuber, 1997) exploit locality as an inductive bias. Many long context Transformers adopt local sparsity as an approximation to dense attention (Child et al., 2019; Beltagy et al., 2020; Zaheer et al., 2020). Locality can be exploited to improve training efficiency via staged training (Press et al., 2021; 2022). The local structure of anticipatory sequences is conducive to the broad class of methods that exploit or depend upon locality. Because anticipation only intervenes to modify the data, it can be seamlessly combined with other modeling innovations, e.g., in the music space: the relative transformer (Huang et al., 2018), the compound word transformer (Hsiao et al., 2021), and RIPO attention (Guo et al., 2022).
@@
Generative music models have not yet reached a broad audience alongside generative models of language (Brown et al., 2020) or images (Ramesh et al., 2022). Slow adoption of music models within creative communities is partly attributable to the difficulty of controlling these models (Briot & Pachet, 2020). Users value agency in human-AI collaborations, preferring to take an active role over more automated solutions (Oh et al., 2018; Roy et al., 2019). We note the proliferation of recent work on music generation (primarily in the audio domain) controlled by text (Forsgren & Martiros, 2022; Kreuk et al., 2023; Agostinelli et al., 2023; Schneider et al., 2023; Huang et al., 2023a; b); we are intrigued by the possibility of applying anticipation to generate symbolic music conditioned on localized text labels (e.g., lyrics). More broadly, we view symbolic music generation and audio music synthesis as complementary, analogous to text generation and speech synthesis. We hope that the Anticipatory Music Transformer may fill a role that is currently underdeveloped in the construction of controllable music generation systems that support the human creative process.
Acknowledgements
This work was done at the Center for Research on Foundation Models (CRFM) at Stanford University. We thank the CRFM and the Stanford Institute for Human-Centered Artificial Intelligence (HAI) for supporting this work. Toyota Research Institute (TRI) and Schmidt Futures provided funds to support this work. The experiments discussed in this paper were conducted on Cloud TPU VMs, provided by the Google TPU Research Cloud (TRC).
Appendix A Ethical Considerations for Generative Models of Music